# Additional dimensions of space and time in the region of deep inelastic processes

### v1

@@ V Conclusions
In this Letter, we have developed the method based on the quantum uncertainty relations and the theory of indirect measurements, which allows to evaluate the module of the group velocity of virtual photons $\gamma^{*}$ and to reveal the type of space-time geometry in the interaction region of high-energy deep inelastic scattering processes. The HERA data show that the speed of virtual photons, $\beta^{*}$, exceeds the speed of light in free space, $\beta^{*}\geq 1$. The superluminal speed means that the virtual photons $\gamma^{*}$ behave tachyon-like. A solution of the problem of the normalization condition $\beta^{*}=1$ at $Q^{2}=0\,\rm{GeV}^{2}$ is proposed in the framework of ideas of ”2T-physics” developed by I. Bars. It is shown that the “admixture” of extra time and extra space dimensions in DIS processes makes it possible to satisfy the normalization condition $\sqrt{3}A_{eff}=1$ as observed in the effective physical (1,3) Minkowski space-time.
The experimental results for virtual photons presented here are based on only a small fraction of the HERA data to illustrate the method, but the analysis can be extended to the full data set, and these investigations are left to future work.
@@ ACKNOWLEDGMENTS
The author thanks M. Wing for reading an earlier version of the manuscript and suggestions, E. Boos, L. Dudko, A. Geiser, N. Nikitin, E. Oborneva, I. Volobuev and other colleagues from SINP MSU and the ZEUS collaboration for discussions and comments. The author is especially grateful to K. G. Gulamov, N. N. Nikolaev and B.S. Yuldashev for their mentorship in the early days. The author expresses his gratitude to the Ministry of Science and Education of the Russian Federation and the DESY Directorate for their long-term financial support and for the hospitality extended to the non-DESY members of the ZEUS collaboration.

### v2

@@ V Conclusions
This communication presents formula (9) for estimating the lower bound of the modulus of the group velocity of a virtual particle. When analysing experimental data, the uncertainties included in the formula should be calculated by methods of indirect measurement theory. To estimate the speed of virtual photons, the combined data from the H1 and ZEUS experiments on deep inelastic $ep$ scattering at the HERA collider were used as input. The HERA data at $Q^{2}\leq$100 GeV2 show that the normalized speed of virtual photons, $\beta^{*}$, exceeds the speed of light in free space, $\beta^{*}\geq 1$. The superluminal speed means that the virtual photons $\gamma^{*}$ behave tachyon-like. This characteristic behavior of virtual photons is consistent with the predictions made by Recami and co-authors in the 1970s and 1980s that tachyons are intermediate states or exchanged objects in elementary processes.
When the normalisation condition $\beta^{*}=1$ is imposed in a real photon limit, the deep relationship between the normalisation parameter $A_{t}$ and the type of spacetime geometry in the interaction domain is revealed. A solution of the problem of the normalisation condition $\beta^{*}=1$ at $Q^{2}=0\,\rm{GeV}^{2}$ is proposed in the framework of ideas of ”Two-time physics” developed by I. Bars. It is shown that by admixing one extra time and one exta space dimensions to 4D spacetime in the domain of DIS processes, it is possible to satisfy the normalisation condition $\sqrt{3}A_{eff}=1$ as observed in the effective physical (1,3) Minkowski spacetime.
For the sake of illustration of the method, the results for virtual photons presented here are based on only a small fraction of the HERA data, but the analysis can be extended to the full data set, and these investigations are left to future work.
@@ ACKNOWLEDGMENTS
The author thanks M. Wing for reading an earlier version of the manuscript and suggestions, E. Boos, L. Dudko, A. Geiser, N. Nikitin, E. Oborneva, I. Volobuev and other colleagues from SINP MSU and the ZEUS collaboration for discussions and comments. The author is especially grateful to K. G. Gulamov, N. N. Nikolaev and B.S. Yuldashev for their mentorship in the early days of this amazing scientific journey. The author expresses his gratitude to the Ministry of Science and Education of the Russian Federation and the DESY Directorate for their long-term financial support and for the hospitality extended to the non-DESY members of the ZEUS collaboration.